OpenStudy (anonymous):

Suppose an equilateral triangle has sides given by 3y - 4 5y - 18 2y + 3 What is the length of each side of the triangle?

OpenStudy (anonymous):

also can you explain how to do this please?

OpenStudy (anonymous):

do you know the definition of equilateral triangle?

OpenStudy (anonymous):

yeah all sides are equal

OpenStudy (anonymous):

so just set the expression for one side equal to another.

OpenStudy (anonymous):

im not getting you bro?? how would i do that? You don't have to give me the answer i jus need you to walk me thru it..

OpenStudy (anonymous):

i am. you have an expression for each side. Side 1 = 3y-4 Side 2 = 5y-18 Side 3 = 2y+3 Set one side equal to another. So set side 1 = side 2 OR side 2= side3 OR side 1 = side 3

OpenStudy (anonymous):

lol im sorry but how would i ''set'' the sides?

OpenStudy (anonymous):

He's saying since it an equalateral triangle all the sides will be the same so: 3y-4=5y-18=2y+3 So the length of the side could be solved using any one of those. 2y+3=0 2y=-3 y=2/3 since a length can never be negative. Do that for all three (the two remaining) equations and you have the lengths of all the sides.

OpenStudy (anonymous):

thank you!